OpenStudy (anonymous):

Find the equation for a parabola with its focus at (0, -3) and a directrix of y = 3. x = y2 x = y2 y = x2 y = x2

OpenStudy (ikram002p):

is this is ur question ?

OpenStudy (tkhunny):

Do you recall the DEFINITION of a parabola, given Focus an dDirectrix? That will help. Do you recall how the Vertex relates to the Focus and Directrix? That will help. Does this parabola open up or down? If it's "down", then the correct answer is not given. And, since two answers are duplicated, I'm guessing you just missed the negative sign on two of them.

OpenStudy (anonymous):

x = -1/12y2 x = 1/12y2 y = 1/12x2 y = -1/12x2

OpenStudy (tkhunny):

Okay, now answer the question. You have all the necessary information if you know only which direction it opens. What is it? Up, Down, Left, or Right?

OpenStudy (anonymous):

up?

OpenStudy (tkhunny):

You're guessing. The focus in inside the parabola. The directrix is outside. Draw them on a graph. Where does the parabola need to be to make these two conditions happen? Answer without a question mark.

OpenStudy (anonymous):

after drawing i got it on the right

OpenStudy (tkhunny):